A patient is to be given 650 ml of a drug over a 5-hour period at...

  1. A patient is to be given 650 ml of a drug over a 5-hour period at 20 drops/ml. What will be the drip rate expressed in gtts/min? Round to the nearest whole number
  1. Your patient has been receiving digoxin (Lanoxin) 125 mcg Q AM. Today his doctor writes a new order: Digoxin 0.25 mg PO Q AM start now. How many 125 mcg tablets will you administer?
  2. At the change of shift you notice 200 ml left to count in the I.V. bag. The I.V. is infusing at 80 mL/hr. How much longer will the I.V. run? (Express you answer in hours and minutes.)
Answer #1

ANSWER.

* Calculation of gtt/min.

       Formula

           gtt/min = Volume(ml) / Time(min) * Drop factor (gtt/ml)

Here

      Volume(ml) = 650 ml

      Time(min) = 300 min (5hr).

      Drop factor = 20 gtt/ml

As per above formula

          gtt/min = 650/300 * 20

                       = 2.17 * 20

                       = 43.4

                       = 43 gtt/min.

* Calculation of number of tablet required.

          Ordered dose = 125 mcg.

          Available dose = 0.25 mg / tablet

    Convert mg to mcg for calculation

                1 mg = 1000 mcg

       So

          Available dose = 0.25 * 1000

                                  = 250 mcg.

Number of tablet required = 125/250

                                             = 0.5 (Half of a tablet) qam.

* Calculation of how much time the IV run.

          Formula

                Rate (ml/hr) = Volume(ml) /Time(hr)

        So

          Time(hr) = Volume(ml) / Rate (ml/hr)

    Here

                      Total volume left in the bag = 200 ml.

                       Rate (ml/hr) = 80 ml/hr.

     As per above formula

                Time = 200 / 80

                          = 2.5 hour

                          = 2 hour and 30 minutes.